OpenStudy (anonymous):

A car accelerates uniformly from +10.0 m/s to +40.0 m/s over a distance of 125 m. How long did it take to go that distance? Show all your work, including the equation used, given and unknown quantities, and any algebra required. Make sure your answer has the correct number of significant figures.

ganeshie8 (ganeshie8):

kinematics ?

OpenStudy (anonymous):

@ganeshie8 yes. please help me.

ganeshie8 (ganeshie8):

first start by putting down whats given

ganeshie8 (ganeshie8):

A car accelerates uniformly from +10.0 m/s to +40.0 m/s over a distance of 125 m initial velocity u = 10 final velocity v = 40 distance s = 125

ganeshie8 (ganeshie8):

t = ?

ganeshie8 (ganeshie8):

which equations we can use ?

ganeshie8 (ganeshie8):

any ideas ?, u have v, u, s and u want t

OpenStudy (anonymous):

isnt initial velocity vi? and final velocity is vf?

OpenStudy (anonymous):

@ganeshie8 vi=10.0, vf=40.0, deltax(distance)=125 t: ? formula: delta d =1/2(vi+vf) delta t 125/ ½ (10.0+40.0) = 5.0 sec

OpenStudy (anonymous):

am i right?

ganeshie8 (ganeshie8):

hey no, we need to use kinematics equations

ganeshie8 (ganeshie8):

v^2-u^2 = 2as

ganeshie8 (ganeshie8):

saw this equation before ? :)

OpenStudy (anonymous):

ohh... um no I havent

OpenStudy (anonymous):

so what is v^2 and u^2 and as?

ganeshie8 (ganeshie8):

a is acceleration s is distance travelled v and u are final/initial velocities

OpenStudy (anonymous):

ok... what about the =2as?

ganeshie8 (ganeshie8):

sorry, thats 2*a*s

ganeshie8 (ganeshie8):

v^2-u^2 = 2as 40^2-10^2 = 2*a*125 solve a first

ganeshie8 (ganeshie8):

once u have a, use below equation to find time, t v = u+at

ganeshie8 (ganeshie8):

give it a try :)

OpenStudy (anonymous):

so it 10.0^2 * 40.0^2= 1700

ganeshie8 (ganeshie8):

v^2-u^2 = 2as 40^2-10^2 = 2*a*125 1600 - 100 = 250a

ganeshie8 (ganeshie8):

solve a, can u ? :)

OpenStudy (anonymous):

@ganeshie8 ummm whats a?

ganeshie8 (ganeshie8):

a = acceleration

OpenStudy (anonymous):

it doest give acceleration on the question

OpenStudy (anonymous):

v^2-u^2 = -1500

ganeshie8 (ganeshie8):

v^2-u^2 = 2as 40^2-10^2 = 2*a*125 1600 - 100 = 250a 1500 = 250a 1500/250 = a 6 = a so, the acceleration a = 6

ganeshie8 (ganeshie8):

Now that we have acceleration a = 6, we can use below equation to find time, t v = u+at

ganeshie8 (ganeshie8):

v = u+at 40 = 10+6t

OpenStudy (anonymous):

oh i thought it would be 10.0^-40.0^2

ganeshie8 (ganeshie8):

solve t

ganeshie8 (ganeshie8):

haaah i see that :) thats why ive put the entire solution, so that it becomes clear to u :)

ganeshie8 (ganeshie8):

u still need to solve t in above equation

OpenStudy (anonymous):

ok 10+6 = 16...

ganeshie8 (ganeshie8):

DONT ignore the 't' attached to 6

ganeshie8 (ganeshie8):

v = u+at 40 = 10+6t subtract 10 both sides 30 = 6t divide 6 both sides 5 = t t = 5

ganeshie8 (ganeshie8):

so, it takes 5 seconds to go to t hat distance of 125 meters

ganeshie8 (ganeshie8):

see if it makes some sense

OpenStudy (anonymous):

I had 5 as my answer earlier...

ganeshie8 (ganeshie8):

good :) 5 = \(\checkmark\)

OpenStudy (anonymous):

why did you say my answer was wrong in the first place.

ganeshie8 (ganeshie8):

oh we both got same answers, weird

ganeshie8 (ganeshie8):

lets see

ganeshie8 (ganeshie8):

Your method :- ``` vi=10.0, vf=40.0, deltax(distance)=125 t: ? formula: delta d =1/2(vi+vf) delta t 125/ ½ (10.0+40.0) = 5.0 sec ```

OpenStudy (anonymous):

yes i got 5

ganeshie8 (ganeshie8):

Yup ! im still thinking why ur method worked

OpenStudy (anonymous):

lol

ganeshie8 (ganeshie8):

-_- have to think, looks like uniform acceleration has something to do wid it

OpenStudy (anonymous):

whats the namw of the formula you gave me? v^2-u^2=2ac

OpenStudy (anonymous):

name*

ganeshie8 (ganeshie8):

thats v^2-u^2=2as

ganeshie8 (ganeshie8):

they're kinematics equations, u may google

OpenStudy (anonymous):

okay i will

ganeshie8 (ganeshie8):

lets get back on this ltr, to see why ur method worked :)

OpenStudy (anonymous):

ok